Calculate the size of the stellar aberration effect in Newtonian physics for
light coming straight down at speed c and a telescope moving straight
sideways at speed v. At what angle θ must we point the telescope to see
the star? Hint: In what direction is the light moving in the reference frame
of the telescope? Express θ in terms of v and c.
1
Expert's answer
2019-10-28T11:28:34-0400

According to the rule of addition of velocities in Newtonian physics, but taking into account the independence of the speed of light from the reference frame "sin\theta = \frac{v}{c}" ,

and since the angle is small, then "\theta = \frac{v}{c}."